OpenStudy (anonymous):

The probability of event A is x, and the probability of event B is y. If the two events are independent, which condition must be true?

OpenStudy (perl):

P(B|A) = P(B) and P(A|B) = P(A)

OpenStudy (anonymous):

P(A|B) = y P(B|A) = x P(A and B) = x + y P(A and B) = xy @perl these are my chioces

OpenStudy (perl):

what is P(A) = ?

OpenStudy (anonymous):

x @perl